# Multivalued (Wess-Zumino-Novikov) Functional in Fluid Mechanics

### v1

@@ Abstract
These equations alone are sufficient for describing a homentropic and barotropic flows, special flows, when the only dynamical variables are components of the particle number 4-current $n^{\mu}=(n^{0},\,n^{0}\bm{v})$, where $n^{0}$ is the particle number density and $\bm{v}$ is the fluid velocity. In this case the continuity equation
@@
A more general, baroclinic flow involves additional dynamical variables. In this case, more equations stem from symmetries other than spacetime diffeomorphisms are necessary. Among them is the gauge symmetry, which yields the continuity equation. In the usual perfect fluid the continuity equation means the particle number conservation (2).
The phenomenon known as the chiral current anomaly presents an obstacle to the conservation of particle number. The issue arises when the Noether current generated by the gauge symmetry, denoted as $I^{\mu}$, is not gauge-invariant; however, its divergence is. In this case, the particle number is not conserved $\partial_{\mu}n^{\mu}\neq 0$, while the equations of motion remain local and gauge-invariant. The chiral anomaly signifies that the flow entrains a reservoir capable supplying and swapping particles.
@@
The deformation of the Hamilton functional is carried out by the multivalued functional introduced by Novikov in 1981 Novikov . Soon after Novikov’s paper, it was recognized that a class of these functionals already appeared in the early work of Wess and Zumino Wess-Zumino , albeit in a coordinate form. Wess and Zumino constructed the functional whose variation replicates the l chiral anomaly effects in the linear response to external gauge field of Weyl fermions. The multivalued functional suitable for hydrodynamics stems from treating the fluid phase space as a manifold of the infinite-dimensional Lie group $\cal G$, in line with Arnold’s approach to hydrodynamics Arnold .
@@
For instance, the canonical stress tensor of a perfect fluid in (1) expressed in canonical variables is
@@
$\pi_{\mu}=p_{\mu}+\partial_{\mu}\psi\,.$ (8)
In a typical fluid, the gauge phase has no physical significance. However, with the chiral anomaly, the scalar $\psi$ takes on a physical meaning. It does not factor into the equation of motion, but enters the Hamilton functional (see Eq.(40) below).
@@
as
$I^{\mu}=j^{\mu}+\tfrac{k}{2}\epsilon^{\mu\nu\lambda\sigma}\partial_{\nu}\psi\,\partial_{\lambda}\pi_{\sigma}\,,$ (10)
where $j^{\mu}$ is the gauge invariant part of the current
@@
The non-gauge-invariant term in (10) represents the part of the ‘electric’ current flowing through the reservoir. As a curl, it drops out of the continuity equation (4), which, when expressed solely in gauge-invariant terms, becomes
@@
The effects of the chiral anomaly are even more apparent when the fluid is charged and exposed to an electromagnetic field. While we do not explore these effects here, we list the equation of motion without providing the derivation. The gauge potential modifies the canonical momentum as $\pi_{\mu}=p_{\mu}+A_{\mu}$, thereby altering the kinetic helicity $\Sigma^{\mu}=\epsilon^{\mu\nu\lambda\sigma}p_{\nu}(\partial_{\lambda}p_{\sigma}+F_{\lambda\sigma})$ by the Maxwell field tensor $F_{\mu\nu}$. Consequently, the equations of motion (1,12) transform into
$\partial_{\mu}\tensor{T}{{}^{\mu}_{\nu}}=F_{\nu\mu}\,j^{\nu},\quad\partial_{\mu}j^{\mu}=\tfrac{k}{4}\,\,\epsilon^{\mu\nu\lambda\sigma}F_{\nu\mu}{}^{\star}F^{\mu\nu}\,,$ (13)
where ${}^{\star}F^{\mu\nu}=\tfrac{1}{2}\epsilon^{\mu\nu\lambda\sigma}F_{\lambda\sigma}$ is the dual tensor. The second relation, which replaces (12), is commonly known as chiral anomaly.
@@
2. Semiclassical Quantization of Fluid Helicity, Vortex Linking and Particle Production.
We begin with a remark on the normalization of the fluid helicity (7) and interpretation of the (yet to obtain) the deformed continuity equation (4).
@@
At the same time, the total helicity is twice the linking number of vortex loops in units of vortex circulation Moreau ; Moffat . Therefore, it is quantized in multiples of the Planck constant as an even number abanov2022axial
@@
[we use the relation $\partial_{\mu}h^{\mu}=\tfrac{k}{4}\,\Omega_{\mu\nu}{}^{\star}\Omega^{\mu\nu}$, where $\Omega_{\mu\nu}=\partial_{\mu}\pi_{\nu}-\partial_{\nu}\pi_{\mu}$ is the 4-vorticity tensor and ${}^{\star}\Omega^{\mu\nu}=\tfrac{1}{2}\epsilon^{\mu\nu\lambda\sigma}\Omega_{\lambda\sigma}$ is the dual tensor]. Integrating this equation over a time interval and across the entire space we find that the LHS of (18) is a change of the total particle number. The integral of the RHS is $k$ times the change of the linking number
@@
Hence, an ‘instanton’ process of changing the linking of vortex loops by $1$ alters the particle number by $k$. Given that the particle number is integer, $k$ is also an integer.
@@
Geometrically, Onsager’s quantization of circulation renders the gauge group (and thereby the entire fluid phase space) compact. The gauge group becomes $U(1)$ making the field $\psi$ to a phase that winds over a circle.
@@
As the particle number is not conserved, the fluid exchanges particles with a reservoir, a medium of massless particles devoid of space-like momentum. We denote the density number of the reservoir constituents by $\overline{n}$, and introduce the density ratio ${S}={\overline{n}}/n$. Subsequently, the fluid energy density, being a function of $n$ should also be a function of ${S}$. This is a setting of a baroclinic flow FootnoteE .
@@
Here $w=\partial_{n}\varepsilon$ is the analog a specific enthalpy, and $u_{\mu}:=n_{\mu}\,/n$ is a relativistic velocity, a 4-unit vector collinear to the particle current. Then we have the relation
@@
Later, we will require the formula for the differential of fluid pressure, designating the fields $\{\psi,\pi,S\}$ as natural variables. In view of (23) we write the standard thermodynamic relation $-P=\varepsilon-nw$ as $-P=p_{\mu}n^{\mu}+\varepsilon$, from which obtain the formula
$-dP-\partial_{S}\varepsilon dS=n^{\mu}dp_{\mu}=$ $n^{\mu}d\pi_{\mu}-n^{\mu}\partial_{\mu}d\psi\,.$ (24)
@@
are carried out by the Lie derivative $\cal L_{\epsilon}$ along a vector field $\epsilon$. The density ratio transform as a scalar
@@
The momenta $\pi_{\mu}$ transform as the covector defined via a form-valued variation. Let $\pi=\pi_{\mu}dx^{\mu}$ is the momentum 1-form. Then
@@
We will employ the covariant Hamiltonian principle, an efficient approach to inviscid fluid dynamics developed by Carter Carter . In this approach, the flow is characterized by the spacetime Hamilton functional $\Lambda[\pi,{S,\psi}]$, which depends on the natural variables: the gauge phase $\psi$, the canonical momentum, and the a scalar field $S$. The Hamiltonian principle asserts that on the equation of motions the Hamilton functional is invariant under the action of the group $\cal G$ Arnold . The action of the group of spacetime diffeomorphisms (27,29) can be seen as variations under physical D’Alembertian displacements of fluid parcels (26). Then the variation of the Hamilton functional vanishes on a physical flow $\delta\Lambda=0$
@@
$\delta_{\epsilon}\Lambda=\int\mathcal{[J^{\mu}}\delta_{\epsilon}\pi_{\mu}+\pi_{S}\mathcal{\delta}_{\epsilon}{S}-I^{\mu}\partial_{\mu}\delta\psi]\,,$ (30)
introduces the conjugate fields: the flow current $\mathcal{J}$, a conjugate to the canonical momentum, the ‘electric’ current $I$, a conjugate to the gauge phase $\psi$, and the conjugate to the density ratio $\pi_{S}$, the time-like ‘momentum’ of reservoir’s particles. Using explicit forms of the variations (27,29) a simple algebra leads to what Carter referred to as the canonical equation of fluid motion FootnoteL
$\mathcal{J}^{\mu}\Omega_{\mu\nu}+\pi_{\nu}(\partial_{\mu}\mathcal{J}^{\mu})=\pi_{S}\,\partial_{\mu}{S}\,.$ (31)
supplemented by the conservation of the electric current (4) followed from (30) under the variation over $\psi$. The first term on the LHS is the force acting on a rotating fluid parcel. It is balanced by the force due to the fluid source (the second term sometime called the ‘ rocket term’) and the ‘heat’ source on the RHS. A notable feature of the canonical equation is that it does not make a reference to a spacetime metric.
The combined result must be gauge-invariant. The gauge phase $\psi$ introduced through the canonical momentum in the second term should not enter the equations. If the flow field $\cal J$ is gauge-invariant, its divergence vanishes and so does the second term. This is the case of the perfect fluid discussed below. However, if $\cal J$ is not gauge-invariant the $\psi$ dependence of the first term must cancel the $\psi$ dependence of the second term. This requirement imposes a nearly prohibiting condition on $\cal\ J$.
@@
$n^{\mu}\Omega_{\mu\lambda}=(\partial_{S}\varepsilon)\partial_{\lambda}{S}\,,$ (33)
which is equivalent to the conservation laws for the perfect fluid given by (1,2,6).
@@
From the canonical equation (33) it follows that the vorticity 2-form $\Omega=\tfrac{1}{2}\Omega_{\mu\nu}({\rm{d}}x^{\mu}\wedge\!{\rm{d}}x^{\nu})$ of the baroclinic flow, where $d{S}\neq 0$, is non-degenerate ${\rm\ det}\,\Omega_{\mu\nu}\neq 0$. Hence, vorticity provides a symplectic structure to the fluid cotangent bundle. [This is in contrast to the homentropic or barotropic flow, where the RHS of (33) vanishes and $n^{\mu}$ is the null-vector of the vorticity tensor].
Then there are four local coordinates $\{\theta,e,\eta,S\}$, in which the symplectic structure (vorticity), takes on a canonical form (Darboux’s theorem): $\Omega=d\theta\wedge de+d\eta\wedge dS$. As a result, the canonical momenta are locally represented by five coordinates and that one of the coordinate is identified with $S$
$\pi=d\psi+\theta de+\eta dS\,,\,$ (34)
We are provided with a map of the 5-dimensional phase space, denoted by $\mathrm{N}^{5}$ to the 4-dimensional spacetime $\mathrm{M}^{4}$: $\mathrm{N}^{5}\to\mathrm{M}^{4}$, where a point of a spacetime $x$ is mapped out from a distinct circle $S^{1}$, represented by the gauge phase $\psi$. The local coordinates of $\mathrm{N}^{5}$ are associated with five Clebsch potentials $\{\psi,\theta,e,\eta,S\}$ Whitham ; Schutz ; Lynden-Bell . Then the physical canonical momentum $\pi=(\partial_{\mu}\psi+\theta\partial_{\mu}e+\eta\partial_{\mu}S)dx^{\mu}$, being the 1-form in $\mathrm{M}^{4}$ could be seen as a push-forward of the 1-form (34) in $\mathrm{N}^{5}$. The map describes a fibration of the phase space $S^{1}\hookrightarrow\mathrm{N}^{5}\to\mathrm{M}^{4}$. The total space ${\mathrm{N}}^{5}$ consists of fibers, with each fiber being a circle $S^{1}$ spanned by the gauge phase one for each point of the spacetime (the base of the bundle). This setup extends the classical Hopf fibration, given by the Hopf map $S^{1}\hookrightarrow S^{3}\to S^{2}$, whose extension to general manifolds is described in Novikov3 . The map is characterized by the invariant, called Hopf-Novikov invariant Khesin2 . The invariant is the integral of the top-form in $\mathrm{N}^{5}$, which is constructed from the pull-back of the canonical momentum Footnote3
@@
The five dimensional phase space allows the following interpretation: our fluid can be seen as a boundary of an auxiliary 5d fluid. We assume that the auxiliary fluid occupies a 5-dimensional half-space $\mathrm{M}_{+}^{5}$ of a closed space $\mathrm{M}^{5}$. The boundary of the half-space is the physical spacetime ${\mathrm{M}}^{4}=\partial\mathrm{M}_{+}^{5}$. We can think about the fifth coordinate as a gauge phase $\psi$. Then the map $\mathrm{M}^{5}\to\mathrm{N}^{5}$ defines the momentum of the auxiliary fluid via (34). Would the 5d fluid occupy the entire space $\mathrm{M}^{5}$ its total helicity would be the invariant $\mathrm{H}$. The helicity of the 5d fluid occupying a half-space is the multivalued Novikov’s functional
@@
modulo the invariant (35) [we set $\Phi$ to vanish at a reference point in the bulk of $\mathrm{M}_{+}^{5}$ and run the integralin (37) from the boundary to this point]. In this sense the functional is multivalued. Consequently, $\Phi$ can not be expressed in a coordinate-free manner. Still, it could be elementary computed in chosen coordinates. Choosing the gauge phase as a fifth coordinate $(\psi,\,x^{\mu})$, the density $\Phi$, modulo an exact 4-form, is
$\Phi\approx\psi(d\pi)^{2}\approx-\partial_{\mu}\psi\,h^{\mu}\,\approx-\partial_{\mu}\psi\,\Sigma^{\mu}.$ (39)
@@
We employ the multivalued functional to deform the Hamilton functional of the perfect fluid
$\Lambda=\Lambda_{0}+\tfrac{k}{2}\Lambda_{WZN}=-\int_{M^{4}}[P+\tfrac{k}{2}\Phi]\,.$ (40)
While the added functional is not uniquely defined, it nonetheless generates local equation of motion (cf.,Novikov2 ). The ambiguity of the functional does not extend to its variation as the invariant $\mathrm{H}$ does not vary.
Unlike $\Lambda_{0}$, the functional is not gauge-invariant. It opens a channel of inflow of the 5d auxiliary fluid into the physical fluid. Nevertheless, the equations of motion maintain gauge invariance.
In the semiclassical fluid, the multivaluedness of $\Lambda$ leads to the quantization of $k$, as we already discussed in the introduction. It follows from the requirement for $e^{(i/\hbar)\Lambda}$ to be single-valued under a global gauge transformation. Consider a transformation $\psi\to\psi+1$, applied at a fixed time. It changes the functional by the total helicity $\Lambda\to\Lambda+\tfrac{k}{2}\int({\bm{p}}\cdot\nabla\times\bm{p})\,d^{3}\bm{x}$. Since the latter is an even integer, $k$ is quantized (cf.,Witten ).
@@
Now we turn to equations of motions generated by the deformed Hamilton functional. We calculate the Noether currents as defined by (30), and subsequently, substitute them into the continuity equation (3) and the canonical equation (31). First, we vary over $\psi$ while holding $\pi$ fixed. With the help of (39) and (24) we obtain the formula for the ‘electric’ current (3) and deformed continuity equation (18).
The next step is to consider a variation of (40) over the canonical momentum. This yields ${\cal\ J^{\mu}}=n^{\mu}+k\epsilon^{\mu\nu\lambda\sigma}\partial_{\nu}\psi\,\partial_{\lambda}\pi_{\sigma}$. The flow field is not gauge-invariant, however, its divergence $\partial_{\mu}\mathcal{J}^{\mu}=\partial_{\mu}n^{\mu}$ is. Now we have all the components to specialize the canonical equation (31). The remaining step is to verify that the gauge phase $\psi$ cancels out: We can see it with the help of the identity
@@
where the ellipsis denotes the cyclic permutation of all five indices, and $z_{\mu}$ is an arbitrary 5-covector. The result is the gauge-invariant canonical form of the Euler equation
$n^{\mu}\Omega_{\mu\nu}+p_{\nu}(\partial_{\mu}n^{\mu})=(\partial_{S}\varepsilon)\,\partial_{\mu}{S}\,$ (42)
which is equivalently represented by the Newtonian form:
$n^{\mu}\partial_{\mu}p_{\nu}+\partial_{\nu}P=-p_{\nu}(\partial_{\mu}n^{\mu})\,,$ (43)
or by the energy-momentum conservation laws (1,6).
@@
Now, it becomes apparent that our fluid is spinning generating a totally antisymmetric spin current $\Sigma_{\mu\nu\lambda}=\frac{1}{2}\epsilon_{\mu\nu\lambda\sigma}\Sigma^{\sigma}$ with helicity being a spin. Given that the vorticity is twice the angular velocity of a rotating fluid parcel, the force in the RHS of (44) represents the spin-orbit coupling. The WZN term gives fluid particles a spin.
@@
11. Advection.
With the help of the continuity equation (18) and the identity (41) we write Eq.(42) as
$(n^{\mu}-k\Sigma^{\mu})\Omega_{\mu\nu}=(\partial_{S}\varepsilon)\,\partial_{\mu}{S}\,.$ (45)
The form of the Euler equation equation is convenient to obtain advection property of the Clebsch potentials. As expected they are the same as in the perfect fluid
$\mathcal{L}\theta=\mathcal{L}e=\mathcal{L}S=0,$ (46)
$\mathcal{L}\eta=\partial_{S}\varepsilon,\quad\mathcal{L}\psi=-nw\,,$ (47)
except that now the Lie derivative $\mathcal{L}=(n^{\mu}-k\Sigma^{\mu})\partial_{\mu}$ acts along the deformed vector field $n^{\mu}-k\Sigma^{\mu}$. Contrary to the perfect fluid this field is not conserved.

### v2

@@ Abstract
These equations alone are sufficient for describing homentropic and barotropic flows, special flows, when the only dynamical variables are components of the particle number 4-current $n^{\mu}=(n^{0},\,n^{0}\bm{v})$, where $n^{0}$ is the particle number density and $\bm{v}$ is the fluid velocity. In this case the continuity equation
@@
A more general, baroclinic flow involves additional dynamical variables. In this case, more equations stem from symmetries other than spacetime diffeomorphisms are necessary. Among them is the gauge symmetry, which yields the continuity equation. In the usual perfect fluid, the continuity equation means the particle number conservation (2).
The phenomenon known as the chiral current anomaly presents an obstacle to the conservation of particle number. The issue arises when the Noether current generated by the gauge symmetry, denoted as $I^{\mu}$, is not gauge-invariant; however, its divergence is. In this case, the particle number is not conserved $\partial_{\mu}n^{\mu}\neq 0$, while the equations of motion remain local and gauge-invariant. The chiral anomaly signifies that the flow entrains a reservoir capable of supplying and swapping particles.
@@
The deformation of the Hamilton functional is carried out by the multivalued functional introduced by Novikov in 1981 Novikov . Soon after Novikov’s paper, it was recognized that a class of these functionals already appeared in the early work of Wess and Zumino Wess-Zumino , albeit in a coordinate form. Wess and Zumino constructed the functional whose variation replicates the l chiral anomaly effects in the linear response to the external gauge field of Weyl fermions. The multivalued functional suitable for hydrodynamics stems from treating the fluid phase space as a manifold of the infinite-dimensional Lie group $\cal G$, in line with Arnold’s approach to hydrodynamics Arnold .
@@
In these terms, the canonical stress tensor of a perfect fluid in (1) expressed in canonical variables is
@@
$\pi_{\mu}=p_{\mu}+\partial_{\mu}\Theta\,.$ (8)
In a typical fluid, the gauge phase has no physical significance. However, with the chiral anomaly, the scalar $\Theta$ takes on a physical meaning. It does not factor into the equation of motion but enters the Hamilton functional (see Eq.(41) below).
@@
as $I^{\mu}=j^{\mu}+\tfrac{k}{2}\epsilon^{\mu\nu\lambda\sigma}\partial_{\nu}\Theta\,\partial_{\lambda}\pi_{\sigma}$, where $j^{\mu}$ is the gauge invariant part of the current
@@
The non-gauge-invariant term in the expression of $I^{\mu}$ can be interpreted as a part of the ‘electric’ current flowing through the reservoir. As a curl, it drops out of the continuity equation (4), which, when expressed solely in gauge-invariant terms, becomes
@@
The effects of the chiral anomaly are even more apparent when the fluid is charged and exposed to an electromagnetic field. Taking an electromagnetic field into account amounts to a modification of the canonical momentum as
$\pi_{\mu}=p_{\mu}+A_{\mu}\,,$ (14)
thereby altering the kinetic helicity
$\Sigma^{\mu}=\epsilon^{\mu\nu\lambda\sigma}p_{\nu}(\partial_{\lambda}p_{\sigma}+F_{\lambda\sigma})$ (15)
by the field tensor $F_{\mu\nu}$. Consequently, the continuity equations (11) transform into
$\partial_{\mu}j^{\mu}=\tfrac{k}{8}\,\,\epsilon^{\mu\nu\lambda\sigma}F_{\mu\nu}F_{\lambda\sigma}\,.$ (16)
This relation is commonly known as chiral anomaly.
We will omit the external gauge field in intermediate formulas in the main text. It can be easily restored using the relation (14).
2. Semiclassical Quantization of Fluid Helicity, Vortex Linking, and Particle Production.
We begin with a remark on the normalization of the fluid helicity (7) and interpretation of the (yet to obtain) deformed continuity equation (4).
@@
At the same time, the total helicity $\int h^{0}d^{3}\bm{x}$ is twice the linking number of vortex loops in units of vortex circulation Moreau ; Moffat . Therefore, it is quantized in multiples of the Planck constant as an even number abanov2022axial
@@
[we use the relation $\partial_{\mu}h^{\mu}=\tfrac{k}{4}\,\Omega_{\mu\nu}{}^{\star}\Omega^{\mu\nu}$, where $\Omega_{\mu\nu}=\partial_{\mu}\pi_{\nu}-\partial_{\nu}\pi_{\mu}$ is the 4-vorticity tensor and ${}^{\star}\Omega^{\mu\nu}=\tfrac{1}{2}\epsilon^{\mu\nu\lambda\sigma}\Omega_{\lambda\sigma}$ is the dual tensor]. Integrating this equation over a time interval and across the entire space we find that the LHS of (19) is a change in the total particle number. The integral of the RHS is $k$ times the change of the linking number
@@
Hence, an ‘instanton’ process of changing the linking of vortex loops by $1$ alters the particle number by $k$. Given that the particle number is an integer, $k$ is also an integer.
@@
Geometrically, Onsager’s quantization of circulation renders the gauge group (and thereby the entire fluid phase space) compact. The gauge group becomes $U(1)$ making the field $\Theta$ into a phase that winds over a circle.
@@
As the particle number is not conserved, the fluid exchanges particles with a reservoir, a medium of massless particles devoid of space-like momentum. We denote the density number of the reservoir constituents by $\overline{n}$ and introduce the density ratio ${S}={\overline{n}}/n$. Subsequently, the fluid energy density, being a function of $n$ should also be a function of ${S}$. This is a setting of a baroclinic flow FootnoteE .
@@
Here $w=\partial_{n}\varepsilon$ is the analog of a specific enthalpy, and $u_{\mu}:=n_{\mu}\,/n$ is a relativistic velocity, a 4-unit vector collinear to the particle current. Then we have the relation
@@
Later, we will require the formula for the differential of fluid pressure, designating the fields $\{\Theta,\pi,S\}$ as natural variables. In view of (24) we write the standard thermodynamic relation $-P=\varepsilon-nw$ as $-P=p_{\mu}n^{\mu}+\varepsilon$, from which obtain the formula
$-dP-\partial_{S}\varepsilon dS=n^{\mu}dp_{\mu}=$ $n^{\mu}d\pi_{\mu}-n^{\mu}\partial_{\mu}d\Theta\,.$ (25)
@@
are carried out by the Lie derivative $\cal L_{\epsilon}$ along a vector field $\epsilon$. The density ratio transforms as a scalar
@@
The momenta $\pi_{\mu}$ transform as the covector defined via a form-valued variation. Let $\pi=\pi_{\mu}dx^{\mu}$ be the momentum 1-form. Then
@@
We will employ the covariant Hamiltonian principle, an efficient approach to inviscid fluid dynamics developed by Carter Carter . In this approach, the flow is characterized by the spacetime Hamilton functional $\Lambda[\pi,\Theta,S]$, which depends on five
$\text{natural variables}:\quad\{\pi,\,\Theta,\,S\},$
the gauge phase $\Theta$, the canonical momentum, and the a scalar field $S$. The Hamiltonian principle asserts that on the equation of motions, the Hamilton functional is invariant under the action of the group $\cal G$ Arnold . The action of the group of spacetime diffeomorphisms (28,30) can be seen as variations under physical D’Alembertian displacements of fluid parcels (27). Then the variation of the Hamilton functional vanishes on a physical flow $\delta\Lambda=0$.
@@
$\delta_{\epsilon}\Lambda=\int\mathcal{[J^{\mu}}\delta_{\epsilon}\pi_{\mu}+\pi_{S}\mathcal{\delta}_{\epsilon}{S}-I^{\mu}\partial_{\mu}\delta\Theta]\,,$ (31)
introduces the conjugate fields: the flow current $\mathcal{J}$, a conjugate to the canonical momentum, the ‘electric’ current $I$, a conjugate to the gauge phase $\Theta$, and the conjugate to the density ratio $\pi_{S}$, the time-like ‘momentum’ of the reservoir’s particles. Using explicit forms of the variations (28,30) a simple algebra leads to what Carter referred to as the canonical equation of fluid motion FootnoteL
$\mathcal{J}^{\mu}\Omega_{\mu\nu}+\pi_{\nu}(\partial_{\mu}\mathcal{J}^{\mu})=\pi_{S}\,\partial_{\nu}{S}\,.$ (32)
supplemented by the conservation of the electric current (4) followed from (31) under the variation over $\Theta$. The first term on the LHS is the force acting on a rotating fluid parcel. It is balanced by the force due to the fluid source (the second term sometime called the ‘ rocket term’) and the ‘heat’ source on the RHS. A notable feature of the canonical equation is that it does not make a reference to a spacetime metric.
The combined result must be gauge-invariant. The gauge phase $\Theta$ introduced through the canonical momentum in the second term should not enter the equations. If the flow field $\cal J$ is gauge-invariant, its divergence vanishes and so does the second term. This is the case of the perfect fluid discussed below. However, if $\cal J$ is not gauge-invariant the $\Theta$ dependence of the first term must cancel the $\Theta$ dependence of the second term. This requirement imposes a nearly prohibiting condition on $\cal\ J$.
@@
$n^{\mu}\Omega_{\mu\nu}=(\partial_{\nu}\varepsilon)_{n}\,,$ (34)
which is equivalent to the conservation laws for the perfect fluid given by (1,2,6). Here $(\partial_{\nu}\varepsilon)_{n}=\partial_{S}\varepsilon\partial_{\nu}S$ is the gradient of energy at a fixed $n$.
@@
From the canonical equation (34) it follows that the vorticity 2-form $\Omega=\tfrac{1}{2}\Omega_{\mu\nu}({\rm{d}}x^{\mu}\wedge{\rm{d}}x^{\nu})$ of the baroclinic flow, where $d{S}\neq 0$, is non-degenerate ${\rm\ det}\,\Omega_{\mu\nu}\neq 0$. Hence, vorticity provides a symplectic structure to the fluid cotangent bundle. [This is in contrast to the homentropic/barotropic flow, where the RHS of (34) vanishes and $n^{\mu}$ is the null-vector of the vorticity tensor].
Then there are four local coordinates $\{\alpha,\,\beta,\,\eta,S\}$, in which the symplectic structure (vorticity), takes on a canonical form (Darboux’s theorem): $\Omega=d\alpha\wedge d\beta+d\eta\wedge dS$. As a result, the canonical momenta are locally represented by five coordinates and that one of the coordinates could be chosen to be $S$
$\pi=d\Theta+\alpha d\beta+\eta dS\,,\,$ (35)
We are provided with a map of the 5-dimensional phase space, denoted by $\mathrm{N}^{5}$ to the 4-dimensional spacetime $\mathrm{M}^{4}$: $\mathrm{N}^{5}\to\mathrm{M}^{4}$, where a point of a spacetime $x$ is mapped out from a distinct circle $\mathrm{S}^{1}$, represented by the gauge phase $\Theta$. The local coordinates of $\mathrm{N}^{5}$ are associated with five Clebsch potentials $\{\Theta,\alpha,\beta,\eta,S\}$ Whitham ; Schutz ; Lynden-Bell . Then the physical canonical momentum $\pi=(\partial_{\mu}\Theta+\alpha\partial_{\mu}\beta+\eta\partial_{\mu}S)dx^{\mu}$, being the 1-form in $\mathrm{M}^{4}$ could be seen as a push-forward of the 1-form (35) in $\mathrm{N}^{5}$. The map describes a fibration of the phase space $\mathrm{S}^{1}\hookrightarrow\mathrm{N}^{5}\to\mathrm{M}^{4}$. The total space ${\mathrm{N}}^{5}$ consists of fibers, with each fiber being a circle $\mathrm{S}^{1}$ spanned by the gauge phase one for each point of the spacetime (the base of the bundle). This setup extends the classical Hopf fibration, given by the Hopf map $\mathrm{S}^{1}\hookrightarrow\mathrm{S}^{3}\to\mathrm{S}^{2}$, whose extension to general manifolds was introduced in Novikov3 . The map is characterized by the invariant, called Hopf-Novikov invariant Khesin2 . The invariant is the integral of the top-form in $\mathrm{N}^{5}$, which is constructed from the pull-back of the canonical momentum Footnote3
@@
The five-dimensional phase space allows the following interpretation: our fluid can be seen as a boundary of an auxiliary 5D fluid. We assume that the auxiliary fluid occupies a 5-dimensional half-space $\mathrm{M}_{+}^{5}$ of a closed space $\mathrm{M}^{5}$. The boundary of the half-space is the physical spacetime ${\mathrm{M}}^{4}=\partial\mathrm{M}_{+}^{5}$. We can think about the fifth coordinate as a gauge phase $\Theta$. Then the map $\mathrm{M}^{5}\to\mathrm{N}^{5}$ defines the momentum of the auxiliary fluid via (35). Would the 5D fluid occupy the entire space $\mathrm{M}^{5}$ its total helicity would be the invariant $\mathrm{H}$. The helicity of the 5D fluid occupying a half-space is the multivalued Novikov’s functional
@@
modulo the invariant (36) [we set $\Phi$ to vanish at a reference point in the bulk of $\mathrm{M}_{+}^{5}$ and run the integral in (38) from the boundary to this point]. In this sense the functional is multivalued. Consequently, $\Phi$ can not be expressed in a coordinate-free manner. Still, it could be elementary computed in chosen coordinates. Choosing the gauge phase as a fifth coordinate $(\Theta,\,x^{\mu})$, the density $\Phi$, modulo an exact 4-form, is
$\Phi=\tfrac{1}{2}\Theta(\Omega\wedge\Omega)\,.$ (40)
@@
We employ the multivalued functional to deform the Hamilton functional of the perfect fluid FootnoteA
$\Lambda=\Lambda_{0}+\tfrac{k}{4}\Lambda_{WZN}=-\int_{M^{4}}P-\tfrac{k}{4}\int_{M^{4}}\Theta(\Omega\wedge\Omega)\,.$ (41)
While the added functional is not uniquely defined, it nonetheless generates a local equation of motion (cf.,Novikov2 ). The ambiguity of the functional does not extend to its variation as the invariant $\mathrm{H}$ does not vary.
Unlike $\Lambda_{0}$, the functional is not gauge-invariant. It opens a channel of inflow of the 5D auxiliary fluid into the physical fluid. Nevertheless, the equations of motion maintain the gauge invariance.
In the semiclassical fluid, the multivaluedness of $\Lambda$ leads to the quantization of $k$, as we already discussed in the introduction. It follows from the requirement for $\exp{(i/\hbar)\Lambda}=\exp{(2\pi i\Lambda)}$ to be single-valued under a global gauge transformation. Consider a transformation $\Theta\to\Theta+1$, applied at a fixed time. It changes the functional by the total helicity $\Lambda\to\Lambda+\tfrac{k}{2}\int({\bm{\pi}}\cdot\nabla\times\bm{\pi})\,d^{3}\bm{x}$. Since the latter is an even integer, $k$ is quantized (cf.,Witten ).
@@
Now we turn to equations of motions generated by the deformed Hamilton functional. We calculate the Noether currents as defined by (31), and subsequently, substitute them into the continuity equation (3) and the canonical equation (32). First, we vary over $\Theta$ while holding $\pi$ fixed. With the help of (40) and (25) we obtain the formula for the ‘electric’ current (3) and deformed continuity equation (19).
The next step is to consider a variation of (41) over the canonical momentum. This yields ${\cal\ J^{\mu}}=n^{\mu}+k\epsilon^{\mu\nu\lambda\sigma}\partial_{\nu}\Theta\,\partial_{\lambda}\pi_{\sigma}$. The flow field is not gauge-invariant, however, its divergence $\partial_{\mu}\mathcal{J}^{\mu}=\partial_{\mu}n^{\mu}$ is. Now we have all the components to specialize the canonical equation (32) and to verify that the gauge phase $\Theta$ cancels out. We can see it with the help of the identity
@@
where the ellipsis denotes the cyclic permutation of all five indices, and $z_{\mu}$ is an arbitrary covector. The result is the canonical form of the Euler equation
$n^{\mu}\Omega_{\mu\nu}+p_{\nu}(\partial_{\mu}n^{\mu})=(\partial_{\nu}\varepsilon)_{n}\,.$ (43)
It’s no surprise that the Euler equation remains unaffected. The WZN term, being independent of metric, does not influence the stress tensor (6).
Expressing the Euler equation in its Newtonian form:
$n^{\mu}\partial_{\mu}p_{\nu}+\partial_{\nu}P=-p_{\nu}(\partial_{\mu}n^{\mu})\,.$ (44)
and using the continuity equation (19) we obtain the Euler equation specified for the chiral anomaly
$(n^{\mu}-k\Sigma^{\mu})\Omega_{\mu\nu}=(\partial_{\nu}\varepsilon)_{n}\,.$ (45)
These equations could be easily extended to include the electromagnetic field.
@@
Now, it becomes apparent that our fluid is spinning generating a totally antisymmetric spin current with a half of helicity $\tfrac{1}{2}\Sigma^{\mu}$ being a spin. Given that the vorticity is twice the angular velocity of a rotating fluid parcel, the force in the RHS of (46) represents the spin-orbit coupling. The WZN term gives fluid particles a spin.
@@
11. Advection and potential vorticity.
he Euler equation along with the continuity equation (19) are the complete set of five equations of motion. The evolution of other physical quantities, such as the density ratio $S$ and the gauge phase $\Theta$, can be readily obtained from (45)
$(n^{\mu}-k\Sigma^{\mu})\partial_{\mu}S=0,\quad(n^{\mu}-k\Sigma^{\mu})\partial_{\mu}\Theta=-nw\,.$
The Clebsch potentials $\alpha$ and $\beta$ are advected similar to $S$, while the ‘thermacy’ $\eta$ obeys the equation $(n^{\mu}-k\Sigma^{\mu})\partial_{\mu}\eta=\partial_{S}\varepsilon$. The Ertel theorem undergoes a similar modification: the potential vorticity $e=(\Sigma^{\lambda}\partial_{\lambda}S)/(nw)$ is conserved $\partial_{\mu}[(n^{\mu}-k\Sigma^{\mu})e]=0$, albeit along a modified vector field $n^{\mu}-k\Sigma^{\mu}$.